OpenStudy (jasonderubis):

A man carries a 10 kg sack of groceries in his arms with a force of 50 N as he walks forward a distance of 10 m. How much work has he done? 0 J 50 J 100 J 500 J

OpenStudy (farcher):

If the question is about the work he does on the sack of potatoes then you have to think about the direction in which the man is moving and the direction in which he is applying the force on the potatoes to keep them at the same height above the ground. Then apply the formula for work done: work done = force x distance moved in the direction of the force.

OpenStudy (anonymous):

500 joules

OpenStudy (anonymous):

Here the mass does not matter as it is acting in a vertical direction Workdone=fs

OpenStudy (anonymous):

@PersianGeek101 I believe that it would end up being 0 J, if we observe the work done on the sack by the man. Drawing a FBD: We see that the displacement of the sack is on the x-axis, but the force being applied to the sack is vertical. If we wish to pursue it further, then mathematically we have: \[\huge \text{W}=\text{F} \cdot \Delta x=\text{F} \Delta x \cos(\theta)\]Since the displacement is perpendicular to the force being applied, then a 90deg angle is formed. cos(90)=0, therefore, the work done by the man on the sack is also 0 J. This would be different if the man were lifting or changing the vertical displacement of the sack.